What is the ratio of two segments where the length of one is 8 centimeters while that of the other is 40 centimeters?
Answer by Theo(13342)

let x = the shorter segment
let y = the longer segment.
the ratio formula says that x/y = 8/40
simplify this to get x/y = 1/5

that's the ratio of x/y.
the ratio of y/x = 40/8 = 5/1

x/y is the same as x:y which means the ratio of x to y.
y/x is the same as y:x which means the ratio of y to x.